OpenStudy (anonymous):

1+x over 3 = 2 Please show steps!

OpenStudy (anonymous):

\[\frac{ 1 + x }{ 3 }= 2? \] is that the question?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

objective is to find x right? so you would need to get the fraction out of the way first and by doing that you would multiply 3 to both sides. which would eliminate the denomenator from the left sides of the equation and that leaves this.. 1 + x = 2 * 3

OpenStudy (anonymous):

and then you would subtract 1 from both sides which would leave x by itself and it would look like this x = (2 * 3) + 1

OpenStudy (anonymous):

and you will find the solution to x! :)

OpenStudy (anonymous):

x=7?